(1 point) Consider the continuously compounded yield curve y(T-0.035-0.0 15є-0.5T Consider a 2-year $ 2500 bond thats redeemable at par and pays semi-annual coupons at a rate of C(2) 596. () Determine the bonds purchase price. Purchase Price $ (i) Determine the duration of the bond to 3 decimals. Duration years Note: Use the purchase price to the closest cent in your duration calculation
Answer #1

(I) The solution is

Bond Payouts Time, t Y(t) 1+Y(t) Discount Factor Present Value
(1) = Coupon Rate * Bond Price + Bond Price (at t=2) (2) (3) =
0.035-0.015*EXP(-0.5*(2))
(4) = 1+(3) (5) = 1/(4) (6) = (1) * (5)
0 0 0.02 1.020 0.980392157 0.00
125 0.5 0.02 1.023 0.977213351 122.15
125 1 0.03 1.026 0.974751936 121.84
125 1.5 0.03 1.028 0.972843557 121.61
2625 2 0.03 1.029 0.971362478 2549.83
SUM 2915.43

Therefore, the purchase price is 2915.43

(II)

Time, t Present Value Weighted Present Value
(1) (2) (from table above) (3) = (1) * (2)
0 0.00 0.00
0.5 122.15 61.08
1 121.84 121.84
1.5 121.61 182.41
2 2549.83 5099.65
SUM 2915.43 5464.98

Duration = Weighted Price Value / Current Market Value = 5464.98/2915.43 = 1.875